OpenStudy (anonymous):

Frank has devised a formula for his catering business that calculates the number of meatballs he needs to prepare. The formula is m=4a+2c, where c= number of children; m= number of meatballs; a= number of adults. How many meatballs are required for a party of 20 adults and 8 children.

OpenStudy (anonymous):

c= number of children = 8; a= number of adults = 20; m=4a+2c; now put the values of a and c; \[m= ( 4 \times 20) + ( 2 \times 8 )= 80+16 = 96 \] m=96; it means you require 96 meatballs for a party of 20 adults and 8 children.

OpenStudy (cj49):

m=4(20)+2(8) m=80+16 m=96

OpenStudy (anonymous):

Thank you, both. (:

OpenStudy (anonymous):

96 is correct